# references.bib

### v1

@@ 4.3 Limitations
Our experimental results provide indication that continuous action masking with zonotopes can improve both the sample efficiency and the final policy of PPO. While the sample efficiency is higher in our experiments, computing the relevant action set adds additional computational load as apparent from the increased computation times (see Appendix A.6). Thus, in practice a tight relevant action set might increase the computation time by more than the computation time for additional samples with standard RL algorithms. Yet, if the relevant action set provides guarantees, e.g., is a set of verified safe actions, this increased computation time might be acceptable.
@@
PPO is the most common RL algorithm. However, off-policy algorithms such as Twin Delayed DPPG (TD3) [Fujimoto2018] and Soft Actor-Critic (SAC) [haarnoja2018] are frequently employed as well. The ray mask and generator mask are conceptually applicable for deterministic policies as used in TD3. Yet, the implications on the gradient must be derived for each RL algorithm and are subject to future work. For the distributional mask, treating the integral in (15) as constant w.r.t. to $\theta$ is a substantial simplification, which might be an explanation of the slightly worse convergence of the distributional mask. To address this in future work, one could approximate the integral with a neural network, which has the advantage that it is easily differentiable.
We focus this work on the integration of a convex relevant action set into RL and assume that an appropriate relevant action set can be obtained. Yet, obtaining this can be a major challenge in practice. On the one hand, the larger the relevant action set is with respect to the action space, the smaller the sample efficiency gain from action masking might get. On the other hand, an accurate relevant action set might require significant computation time to obtain and could lead to a high-dimensional generator space, which could be challenging for the generator mask. Thus, future work should investigate efficient methods to obtain sufficiently tight relevant action sets.
@@ 5 Conclusion
We propose action masking methods for continuous action spaces that focus the exploration on the relevant part of the action set. In particular, we extend previous work on continuous action masking to general convex relevant action sets. To this end, we introduce three masking methods and derive their implications on the gradient of PPO. We empirically evaluate the methods on three benchmarks and observe that the generator mask and ray mask perform best. If the relevant action set can be described by a zonotope with fixed generator dimensions and the policy follows a normal distribution, the generator mask is straightforward to implement. If the assumptions for the generator mask cannot be fulfilled, the ray mask is recommended based on our experiments. Because of the subpar performance and longer computation time of the distributional mask, it needs to be further improved. Future work should also investigate a broad range of benchmarks to more clearly identify the applicability and limits of continuous action masking with expressive convex sets.
@@ Appendix A Appendix
A.1 Linear transformation of multivariate normal distributions
@@
Applying the linear mapping of a Zonotope $\langle c,G\rangle_{\mathcal{Z}}$ for the multivariate normal distribution $a\sim\mathcal{N}(\mu,\Sigma)$ results in the multivariate normal distribution $b=c+Ga\sim\mathcal{N}(c+G\mu,G\Sigma G^{T}$).
@@
The moment-generating function of the random variable $a$ is $M_{a}(t)=\mathbb{E}_{a}\left[\exp\left(t^{T}a\right)\right]$, and for $b$ it is
$\begin{split}M_{b}(t)&=\mathbb{E}_{a}\left[\exp\left(t^{T}(c+Ga)\right)\right]\\ &=\exp\left(t^{T}c\right)\mathbb{E}_{a}\left[\exp\left(t^{T}Ga\right)\right]\\ &=\exp\left(t^{T}c\right)M_{a}(G^{T}t).\end{split}$ (21)
Since $a$ is normally distributed, we can write $M_{a}(t)=\exp\left(t^{T}\mu+0.5t^{T}\Sigma t\right)$, and hence
$\begin{split}M_{b}(t)&=\exp\left(t^{T}c\right)\exp\left(t^{T}G\mu+0.5t^{T}G\Sigma G^{T}t\right)\\ &=\exp\left(t^{T}(G\mu+c)+0.5t^{T}G\Sigma G^{T}t\right),\end{split}$ (22)
which is the moment-generating function for the variable b $\sim\mathcal{N}(c+G\mu,G\Sigma G^{T})$. ∎
@@ A.2 Policy gradient for the generator mask
We can state the following, for the special case, when the inverse of $G$ exists.
@@ A.3.1 General case
Given an initial state $s_{0}$, an input trajectory $a$, and a disturbance trajectory $w$, we denote the solution of (2) at time $t$ as $\xi_{t}(s_{0},a,w)$. The reachable set of (2) at some time $t$ given an uncertain initial state $\mathcal{S}_{0}$ and an input trajectory $a$ is denoted as
$\mathcal{R}^{e}_{t}(\mathcal{S}_{0},a)=\big{\{}\xi_{t}(s_{0},a,w)\colon\exists s_{0}\in\mathcal{S}_{0},\exists w\in\mathcal{W}\big{\}}.$ (28)
The reachable set over the time interval $[0,t]$ is defined as
[$\mathcal{R}^{e}_{[0,t]}(\mathcal{S}_{0},a)=\bigcup_{\tau\in[0,t]}\mathcal{R}^{e}_{\tau}(\mathcal{S}_{0},a).$ (29)
For a single time step, we can compute the reachable set given a set of inputs $\tilde{\mathcal{A}}\subseteq\mathcal{A}$ as
$\mathcal{R}^{e}_{\Delta t}(\mathcal{S}_{0},\tilde{\mathcal{A}})=\big{\{}\xi_{t}(s_{0},a,w)\colon\exists x_{0}\in\mathcal{S}_{0},\exists a\in\tilde{\mathcal{A}},\exists w\in\mathcal{W}\big{\}}.$ (30)
It can be prohibitively hard to compute the reachable set exactly, which is why we generally compute overapproximations $\mathcal{R}(\cdot)\supseteq\mathcal{R}^{e}(\cdot)$.
To guarantee constraint satisfaction over an infinite time horizon, we use a relevant state set $\mathcal{S}^{r}$. This set is defined such that constraint satisfaction is guaranteed for all time as long as the system state remains inside it. Formally, it is defined such that
$\forall\mathcal{S}\subseteq\mathcal{S}^{r},\exists a\in\mathcal{A}\colon\mathcal{R}_{\Delta t}(\mathcal{S},a)\subseteq\mathcal{S}^{r},\mathcal{R}_{[0,\Delta t]}(\mathcal{S},a)\subseteq\mathcal{S}.$ (31)
Such sets are commonly called control invariant. For an efficient computation, the reader is referred to [schaefer2024scalable].
We compute the relevant action set as the largest set of inputs that allows us to keep the system in the relevant state set in the next step. We define a parameterized action set $\mathcal{A}^{r}(p)$, where $p\in\mathbb{R}^{n^{p}}$ is a parameter vector. The relevant action set at some time step $k$ is then computed with the optimal program
$\!\max_{p}$ $\text{Vol}\left(\mathcal{A}^{r}_{k}(p)\right)$ (32)
subject to $\mathcal{A}^{r}_{k}(p)\subseteq\mathcal{A}$
$\mathcal{R}_{k+\Delta t}(\mathcal{S}_{k},\mathcal{A}^{r}_{k}(p))\subseteq\mathcal{S}^{r}$
$\mathcal{R}_{[k,k+\Delta t]}(\mathcal{S}_{k},\mathcal{A}^{r}_{k}(p))\subseteq\mathcal{S},$
where $\texttt{Vol}(\cdot)$ is a function representing the set volume. In the following, we provide a detailed formulation of (32) as an exponential cone program.
@@ A.3.2 Exponential cone program
With regard to the input set parameterization, we consider a template zonotope with a predefined template generator matrix $\tilde{G}$ and variable generator scaling factors. We then use the vector $p^{G}\in\mathbb{R}^{n^{g}}_{>0}$ of generator scaling factors to define the generator matrix. The parameterized template zonotope is then given by
$\mathcal{A}^{r}(p)=c+\tilde{G}\,diag(p^{G}),$ (34)
where $p=\begin{bmatrix}c&p^{G}\end{bmatrix}^{\top}$.
We denote the center and generator matrix of a zonotope $\mathcal{Z}$ by $c^{\mathcal{Z}}$ and $G^{\mathcal{Z}}$, respectively. Based on the linearized system dynamics, and using $\mathcal{R}$ as a shorthand for $\mathcal{R}_{k+\Delta t}(\cdot)$, we have $G^{\mathcal{R}}=\begin{bmatrix}AG^{\mathcal{S}_{0}}&BG^{\mathcal{A}^{r}_{k}}&G^{\mathcal{W}}\end{bmatrix}$ and $c^{\mathcal{R}}=\begin{bmatrix}Ac^{\mathcal{S}_{0}}&Bc^{\mathcal{A}^{r}_{k}}&c^{\mathcal{W}}\end{bmatrix}$.
We follow the approach in [sadraddini2019linear] to formulate the set containment constraints in (32) as linear constraints. Since we already consider trajectory curvature in the disturbance, we only need to guarantee that the relevant input set is contained in the feasible input set and that the reachable set of the next time step is contained in the relevant state set. We define the auxiliary variables $\Gamma^{\mathcal{R}}\in\mathbb{R}^{n^{\mathcal{S}^{r}}\times n^{\mathcal{R}}}$, $\beta^{\mathcal{R}}\in\mathbb{R}^{n^{\mathcal{S}^{r}}}$, $\Gamma^{\mathcal{A}^{r}_{k}}\in\mathbb{R}^{n^{\mathcal{A}}\times n^{\mathcal{A}^{r}_{k}}}$, and $\beta^{\mathcal{A}^{r}_{k}}\in\mathbb{R}^{n^{\mathcal{A}}}$ and solve
$\!\max_{p}$ $\text{Vol}\left(\mathcal{A}^{r}_{k}(p)\right)$ (35)
@@
$G^{\mathcal{A}^{r}_{k}}-G^{\mathcal{A}}\Gamma^{\mathcal{A}^{r}_{k}}=\mathbf{0}$
$c^{\mathcal{A}}-c^{\mathcal{A}^{r}_{k}}-G^{\mathcal{A}}\beta^{\mathcal{A}^{r}_{k}}=\mathbf{0}$
$\|\begin{bmatrix}\Gamma^{\mathcal{A}^{r}_{k}}&\beta^{\mathcal{A}^{r}_{k}}\end{bmatrix}\|_{\infty}\leq 1,$
where $\mathbf{0}$ are zero matrices of appropriate dimensions. By vectorizing $\Gamma^{\mathcal{R}}$ and $\Gamma^{\mathcal{A}^{r}_{k}}$ and resolving the absolute value constraints, we can obtain a formulation with purely linear constraints. Since computing the exact volume of a zonotope is computationally expensive, we approximate it by computing the geometric mean of the parameterization vector:
$\texttt{Vol}(\mathcal{A}^{r}_{k}(p))=\left(\prod_{i}p_{i}\right)^{\frac{1}{n^{p}}}.$ (36)
With the cost function from (36), the problem in (35) represents an exponential cone program, which is convex and can be efficiently computed [boyd2007tutorial].

### v2

@@ 5.3 Discussion and limitations
Our experimental results indicate that continuous action masking with zonotopes can improve both the sample efficiency and the final policy of PPO. While the sample efficiency is higher in our experiments, computing the relevant action set adds computational load as shown by the increased computation times (see Appendix A.6). Thus, in practice, a tight relevant action set might require more computation time than the additional samples for standard RL algorithms. Yet, if the relevant action set provides guarantees, e.g., is a set of verified safe actions, this increased computation time is often acceptable. Additionally, the computational effort for the masks differs. Given a relevant action zonotope, the generator mask adds a matrix-vector multiplication, which scales quadratically with action dimensions, the ray mask is dominated by the computation of the boundary points, which scales polynomially with action dimensions [kulmburg2021co], and the distributional mask scales cubically with the dimension of the action space due to the mixing time of the hit-and-run algorithm [lovasz_hit-and-run_2006]. Note that for the Walker2D environment with six action dimensions, the computation time for the hit-and-run algorithm is so high that the distributional mask evaluation was infeasible.
The ray mask and generator mask are based on functions $g_{\text{R}}$ and $g_{\text{G}}$ that map to relevant actions. There are two different approaches of incorporating these functions into RL algorithms. One is to apply the mapping as part of the environment on an action that is sampled by a standard RL policy. The second option, which we use, is to consider the masking as part of the policy, which creates the relevant policy as in (5). This has three main benefits over integrating masking as part of the environment. First, the actions passed to the environment are better interpretable, e.g., for the generator mask, adding the masking mapping function to the environment leads to a policy that samples actions from the generator space, which commonly does not have an intuitive interpretation. Second, more formulations of the functional $h$ are possible, e.g., the distributional mask, and, third, the mapping function can be included in the gradient calculation. For mathematically sound masking as part of the policy, the correct gradient needs to be derived for each RL algorithm, and the standard RL algorithms need to be adapted accordingly. However, the empirical benefit could be minor. Thus, future work should investigate the significance of the correct gradient on a variety of tasks. Note that we showed in Proposition 1 that for the ray mask, the PPO gradient with respect to the original policy and relevant policy is the same. Thus, for the ray mask, it does not matter if it is viewed as part of the policy or the environment.
@@
PPO is a common RL algorithm. However, off-policy algorithms such as Twin Delayed DDPG (TD3) [Fujimoto2018] and Soft Actor-Critic (SAC) [haarnoja2018] are frequently employed as well. The ray mask and generator mask are conceptually applicable for deterministic policies as used in TD3. Yet, the implications on the gradient must be derived for each RL algorithm and are subject to future work. For the distributional mask, treating the integral in (15) as constant with respect to $\theta$ is a substantial simplification, which might be an explanation for the slightly worse convergence of the distributional mask since this introduces an off-policy bias. To address this in future work, one could approximate the integral with a neural network, which has the advantage that it is easily differentiable.
We focus this work on the integration of a convex relevant action set into RL and assume that an appropriate relevant action set can be obtained. While convex sets are a significant generalization from previous work [Krasowski2023b.ProvablySafeRLSurvey], they might not be sufficient for some applications, e.g., tasks where relevant action sets are disjoint. Thus, future work could include investigating hybrid RL approaches [neunert2020continuous] to increase the applicability to multiple convex sets or non-convex sets, such as constrained polynomial zonotopes [Kochdumper2023_constrainedzono]. Further, obtaining the relevant action set can be a major challenge in practice, in particular when the relevant action set is different for each state. Such a high state-dependency is likely when the notion of relevance is safety, while for other definitions of action relevance, the relevant action set might be easy to pre-compute, e.g., excluding high steering angles at high velocities. Additionally, there might be an optimal precision of the relevant action set due to two opposing mechanisms. On the one hand, the larger the relevant action set is with respect to the action space, the smaller the sample efficiency gain from action masking might get. On the other hand, a tight relevant action set might require significant computation time to obtain. Thus, future work should investigate efficient methods to obtain sufficiently tight relevant action sets.
@@ 6 Conclusion
We propose action masking methods for continuous action spaces that focus the exploration on the relevant part of the action set. In particular, we extend previous work on continuous action masking from using intervals as relevant action sets to using convex sets. To this end, we have introduced three masking methods and have derived their implications on the gradient of PPO. We empirically evaluated our methods on four benchmarks and observed that the generator mask and ray mask perform best. If the relevant action set can be described by a zonotope with fixed generator dimensions and the policy follows a normal distribution, the generator mask is straightforward to implement. If the assumptions for the generator mask cannot be fulfilled, the ray mask is recommended based on our experiments. Because of subpar performance and longer computation time, the distributional mask needs to be further improved. Future work should also investigate a broad range of benchmarks to identify the applicability and limits of continuous action masking with convex sets more clearly.
Acknowledgments and Disclosure of Funding
We thank Matthias Killer for conducting preliminary experiments. We gratefully acknowledge that this project was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – SFB 1608 – 501798263, AL 1185/9-1, AL 1185/33-1, and the Bavarian Research Foundation project STROM (Energy - Sector coupling and microgrids).
@@ Appendix A Appendix
A.1 Proof of bijectivity of the ray mapping function.
@@
The mapping function $g:\mathcal{A}\rightarrow\mathcal{A}^{r}$ of the ray mask $g(a)=c+\frac{\lambda_{\mathcal{A}^{r}}(a)}{\lambda_{\mathcal{A}}(a)}(a-c)$ is bijective.
@@
We show that $g(a)$ is bijective by showing that the function is both injective and surjective.
For any convex set $\mathcal{A}^{r}$ with center $c$, we can construct a ray from $c$ through any $a^{r}\in\mathcal{A}^{r}$ as $r_{\gamma}(t)=c+\gamma t$, where $\gamma=\frac{a-c}{\|a-c\|_{2}}$, and $t\in\left[0,t_{max}\right]$. Since $\mathcal{A}^{r}\subseteq\mathcal{A}$, this also holds $\forall a\in\mathcal{A}$. By construction, any two distinct rays only intersect in $c$. For all points on a given ray, the scaling factors $\lambda_{\mathcal{A}}$ and $\lambda_{\mathcal{A}^{r}}$ are constants, allowing us to rewrite $g(a)$ for all $a=r_{\gamma}(t)$ as the linear function:
$g\left(r_{\gamma}(t)\right)=c+\frac{\lambda_{\mathcal{A}^{r}}}{\lambda_{\mathcal{A}}}\left(r_{\gamma}(t)-c\right).$
To show that $g(a)$ is injective, i.e. $\forall a_{1},a_{2}\in\mathcal{A}$, if $a_{1}\neq a_{2}\implies g(a_{1})\neq g(a_{2})$, consider the following two cases for $a_{1}\neq a_{2}$.
Case 1: If $a_{1}$ and $a_{2}$ are on the same ray, then, $a_{1}=r_{\gamma}(t_{1})$ and $a_{2}=r_{\gamma}(t_{2})$ with $t_{1}\neq t_{2}$. Since $g\left(r_{\gamma}(t)\right)$ is linear and thereby monotonic, it follows that $g\left(r_{\gamma}(t_{1})\right)\neq g\left(r_{\gamma}(t_{2})\right)$ and consequently $g(a_{1})\neq g(a_{2})$.
Case 2: Otherwise, $a_{1}$ and $a_{2}$ are on different rays and $a_{1}\neq a_{2}\neq c$, so $g(a_{1})\neq g(a_{2})$ follows directly as the rays only intersect in $c$.
Thus, $g(a)$ is injective.
For showing $g(a)$ to be surjective, it is sufficient to show that $\forall a^{r}\in\mathcal{A}^{r}$, there exists an $a\in\mathcal{A}$, for which $g(a)=a^{r}$.
Consider the ray $r_{\gamma}(t)$ passing through $a^{r}$. We know that $g\left(r_{\gamma}(0)\right)=g(c)=c$, and $g\left(r_{\gamma}(t_{max})\right)$ is the boundary point of $\mathcal{A}^{r}$ from $c$ in the direction $d$. Moreover, $a^{r}$ lies on the line segment between $c$ and $g\left(r_{\gamma}(t_{\max})\right)$. Since $g\left(r_{\gamma}(t)\right)$ is linear and continuous, according to the intermediate value theorem, $\exists t^{*}\in\left[0,t_{\max}\right]$, for which $g\left(r_{\gamma}(t^{*})\right)=a^{r}$ for any $a^{r}\in\mathcal{A}^{r}$ along the ray $r_{\gamma}(t)$. As we can construct such a ray through any point in $\mathcal{A}^{r}$, we have shown that for every $a^{r}\in\mathcal{A}^{r}$, there exists an $a\in\mathcal{A}$ such that $g(a)=a^{r}$, thus proving surjectivity. ∎
@@ A.2 Policy gradient for the generator mask
We can state the following for the special case when the inverse of $G$ exists.
@@ A.3.1 General case
Given an initial state $s_{0}$, an input trajectory $a_{(\cdot)}$, and a disturbance trajectory $w_{(\cdot)}$, we denote the solution of (2) at time $t$ as $\xi_{t}(s_{0},a_{(\cdot)},w_{(\cdot)})$. Assuming a sampled controller with piecewise-constant input $a$, we define the reachable set of (2) after one time step $\Delta t$ given a set of initial states $\mathcal{S}_{0}$ and a set of possible inputs $\tilde{\mathcal{A}}\subseteq\mathcal{A}$ as
$\mathcal{R}^{e}_{\Delta t}(\mathcal{S}_{0},\tilde{\mathcal{A}})=\big{\{}\xi_{\Delta t}(s_{0},a,w_{(\cdot)})\,\big{|}\,\exists x_{0}\in\mathcal{S}_{0},\exists a\in\tilde{\mathcal{A}},\forall t\colon\exists w_{t}\in\mathcal{W}\big{\}}.$ (26)
The reachable set over the time interval $[0,\Delta t]$ is defined as
[$\mathcal{R}^{e}_{[0,\Delta t]}(\mathcal{S}_{0},\tilde{\mathcal{A}})=\bigcup_{\tau\in[0,\Delta t]}\mathcal{R}^{e}_{\tau}(\mathcal{S}_{0},\tilde{\mathcal{A}}).$ (27)
For many system classes it is impossible to compute the reachable set exactly [platzer2007hybrid], which is why we generally compute overapproximations $\mathcal{R}(\cdot)\supseteq\mathcal{R}^{e}(\cdot)$.
To guarantee constraint satisfaction over an infinite time horizon, we use a relevant state set $\mathcal{S}^{r}$. In this work, we choose $\mathcal{S}^{r}$ to be a robust control invariant set in the sense that we can always find an input which guarantees that the reachable set at the next time step is contained in $\mathcal{S}^{r}$ and that all state constraints are satisfied in the time interval in between [schaefer2024scalable]:
$\exists a\in\mathcal{A}\colon\mathcal{R}_{\Delta t}(\mathcal{S}^{r},a)\subseteq\mathcal{S}^{r},\mathcal{R}_{[0,\Delta t]}(\mathcal{S}^{r},a)\subseteq\mathcal{S}.$ (28)
We compute the relevant action set as the largest set of inputs that allows us to keep the system in the relevant state set in the next time step. We define a parameterized action set $\mathcal{A}^{r}(p)$, where $p\in\mathbb{R}^{n^{p}}$ is a parameter vector. The relevant action set is then computed with the optimal program
$\!\max_{p}$ $\widetilde{\texttt{Vol}}\left(\mathcal{A}^{r}(p)\right)$ (29)
subject to $\mathcal{A}^{r}(p)\subseteq\mathcal{A}$
$\mathcal{R}_{\Delta t}(\mathcal{S}_{0},\mathcal{A}^{r}(p))\subseteq\mathcal{S}^{r}$
$\mathcal{R}_{[0,\Delta t]}(\mathcal{S}_{0},\mathcal{A}^{r}(p))\subseteq\mathcal{S},$
where $\widetilde{\texttt{Vol}}(\cdot)$ is a proxy function for the volume in the sense that a maximization of $\widetilde{\texttt{Vol}}(\cdot)$ is suboptimally maximizing the volume. In the following, we provide a detailed formulation of (29) as an exponential cone program.
@@ A.3.2 Exponential cone program
With regard to the input set parameterization, we consider a template zonotope with a predefined template generator matrix $\tilde{G}$. We use the vector $\tilde{p}\in\mathbb{R}^{P}_{>0}$ of generator scaling factors to scale the generator matrix. The parameterized template zonotope is then given by
$\mathcal{A}^{r}(p)=\langle c,\tilde{G}\,diag(\tilde{p})\rangle_{\mathcal{Z}},$ (31)
where $p=\begin{bmatrix}c&\tilde{p}\end{bmatrix}^{\top}$. In the following, we denote the center and generator matrix of a specific zonotope $\mathcal{Z}$ by $c^{\mathcal{Z}}$ and $G^{\mathcal{Z}}$, respectively. A zonotope $\mathcal{Z}_{1}$ is contained in a zonotope $\mathcal{Z}_{2}$ if there exist $\Gamma\in\mathbb{R}^{P^{\mathcal{Z}_{2}}\times P^{\mathcal{Z}_{1}}}$, $\beta\in P^{\mathcal{Z}_{2}}$, such that [sadraddini2019linear, Corollary 4]
$G^{\mathcal{Z}_{1}}$ $=G^{\mathcal{Z}_{2}}\Gamma$ (32)
$c^{\mathcal{Z}_{2}}-c^{\mathcal{Z}_{1}}$ $=G^{\mathcal{Z}_{2}}\beta$
$\|[\Gamma,\beta]\|_{\infty}$ $\leq 1.$
Based on the linearized system dynamics in (30), the definitions from (4), and using $\mathcal{R}$ as a shorthand for $\mathcal{R}_{\Delta t}(\cdot)$, we have $G^{\mathcal{R}}=\begin{bmatrix}AG^{\mathcal{S}_{0}}&BG^{\mathcal{A}^{r}}&G^{\mathcal{W}}\end{bmatrix}$ and $c^{\mathcal{R}}=\begin{bmatrix}Ac^{\mathcal{S}_{0}}&Bc^{\mathcal{A}^{r}}&c^{\mathcal{W}}\end{bmatrix}$. Since we already consider trajectory curvature in the disturbance, we only need to guarantee that the relevant input set is contained in the feasible input set and that the reachable set of the next time step is contained in the relevant state set. Using the containment condition from (32), we define the auxiliary variables $\Gamma^{\mathcal{R}}$, $\beta^{\mathcal{R}}$, $\Gamma^{\mathcal{A}^{r}}$, and $\beta^{\mathcal{A}^{r}}$ and solve
$\!\max_{p}$ $\widetilde{\texttt{Vol}}\left(\mathcal{A}^{r}(p)\right)$ (33)
@@
$G^{\mathcal{A}^{r}}-G^{\mathcal{A}}\Gamma^{\mathcal{A}^{r}}=\mathbf{0}$
$c^{\mathcal{A}}-c^{\mathcal{A}^{r}}-G^{\mathcal{A}}\beta^{\mathcal{A}^{r}}=\mathbf{0}$
$\|\begin{bmatrix}\Gamma^{\mathcal{A}^{r}}&\beta^{\mathcal{A}^{r}}\end{bmatrix}\|_{\infty}\leq 1,$
where $\mathbf{0}$ are zero matrices of appropriate dimensions. By vectorizing $\Gamma^{\mathcal{R}}$ and $\Gamma^{\mathcal{A}^{r}}$ and resolving the infinity norm constraints [bertsimas1997introduction, Sec. 1.3], we can obtain a formulation with purely linear constraints. Since computing the exact volume of a zonotope is computationally expensive, we approximate it by computing the geometric mean of the parameterization vector:
$\widetilde{\texttt{Vol}}(\mathcal{A}^{r}(p))=\left(\prod_{i}\tilde{p}_{i}\right)^{\frac{1}{P}}.$ (34)
With the cost function from (34), the problem in (33) is an exponential cone program, which is convex and can be efficiently computed [boyd2007tutorial, Sec. 2.5].